# Active Adaptive Experimental Design for Treatment Effect Estimation with Covariate Choices

### v1

@@ Abstract
This study designs an adaptive experiment for efficiently estimating average treatment effect (ATEs). We consider an adaptive experiment where an experimenter sequentially samples an experimental unit from a covariate density decided by the experimenter and assigns a treatment. After assigning a treatment, the experimenter observes the corresponding outcome immediately. At the end of the experiment, the experimenter estimates an ATE using gathered samples. The objective of the experimenter is to estimate the ATE with a smaller asymptotic variance. Existing studies have designed experiments that adaptively optimize the propensity score (treatment-assignment probability). As a generalization of such an approach, we propose a framework under which an experimenter optimizes the covariate density, as well as the propensity score, and find that optimizing both covariate density and propensity score reduces the asymptotic variance more than optimizing only the propensity score. Based on this idea, in each round of our experiment, the experimenter optimizes the covariate density and propensity score based on past observations. To design an adaptive experiment, we first derive the efficient covariate density and propensity score that minimizes the semiparametric efficiency bound, a lower bound for the asymptotic variance given a fixed covariate density and a fixed propensity score. Next, we design an adaptive experiment using the efficient covariate density and propensity score sequentially estimated during the experiment. Lastly, we propose an ATE estimator whose asymptotic variance aligns with the minimized semiparametric efficiency bound.
@@ 1 Introduction
Experimental approaches play a pivotal role in uncovering causality in various fields of science and industrial applications, such as epidemiology and economics Chow SC (2005); van der Laan (2008); Hahn et al. (2011); Chow & Chang (2011); FDA (2019). This study focuses on cases with binary treatments. Typically, one of these treatments corresponds to the treatment, and the other to the control Imbens & Rubin (2015). Among the quantities representing causal effects, this study focuses on an average treatment effect (ATE), which defines a causal effect as the difference between the expected outcomes of binary treatments.
Treatments are an abstraction of an experimenter’s choices, including drugs, online advertisements, and economic policies. Depending on the literature, treatments are referred to differently, such as actions and arms. A difference in outcomes of binary treatments (treatment effect) is often considered a causal effect (Imbens & Rubin, 2015). However, due to the counterfactual property, individual treatment effects (ITEs) cannot be observed. Therefore, we use ATEs, expected values of ITEs, to capture causality.
A fundamental experimental approach for estimating ATEs is a randomized control trial (RCT). In an RCT, one of the binary treatments is randomly assigned to each experimental unit to obtain an unbiased estimator of ATEs. While RCTs are the gold standard and reliable (Hariton & Locascio, 2018), they often require large sample sizes, which can be costly. To reduce the sample size as possible, or to reduce an estimation error given a sample size, adaptive experimental designs have gained increasing attention across various fields van der Laan (2008); Hahn et al. (2011); Chow & Chang (2011); Villar et al. (2015); FDA (2019).
For estimating ATE estimation efficiently (smaller asymptotic variances), van der Laan (2008) and Hahn et al. (2011) design adaptive experiments that optimize the propensity score (treatment-assignment probability) [^1] The propensity score is a term used mainly in observational studies, but we use this term in experimental studies for simplicity.. They show that optimizing the propensity score can minimize an asymptotic variance of ATE estimators. Note that an asymptotic variance can be interpreted as an asymptotic mean squared error (MSE) of ATE estimation. Furthermore, in statistical hypothesis testing, as an asymptotic variance decreases, the sample size required for hypothesis testing also decreases. Tabord-Meehan (2022) and Kato et al. (2020) design refined experiments based on this approach.
@@
Firstly, we derive the semiparametric efficiency bound, a lower bound of the asymptotic variance, which is a functional of a covariate density and a propensity score. Based on this efficiency bound, we compute the efficient covariate density and propensity score that minimize the lower bound. Then, we design an adaptive experiment that sequentially estimates the efficient covariate density and propensity score. In the proposed experiment, an experimental unit is sampled following an adaptively estimated efficient covariate density and randomly assigned treatments following an estimated efficient propensity score. At the experiment’s end, the experimenter obtains an estimator whose asymptotic variance aligns with the semiparametric efficiency bound. We employ the Augmented Inverse Probability Weighting (AIPW) estimator with covariate shift adaptation using the Importance Weighting (AIPWIW). We refer to the setting of an adaptive experiment as active adaptive experimental design with covariate choice and our designed experiment as the Active-Adaptive-Sampling (AAS)-AIPWIW experiment.
@@
Our method is inspired by three lines of work: adaptive experimental design for efficient ATE estimation (van der Laan, 2008; Hahn et al., 2011), active learning using a covariate shift (Sugiyama, 2006), and off-policy evaluation under a covariate shift (importance sampling) (Uehara et al., 2020). van der Laan (2008), Hahn et al. (2011), Tabord-Meehan (2022), and Kato et al. (2020) design adaptive experiments for estimating ATEs with smaller asymptotic variances by optimizing propensity scores. Our result generalizes their work by incorporating the covariate density optimization. In active learning of regression models, Sugiyama (2006) demonstrates that the asymptotic variance of regression coefficient estimators can be reduced by appropriately shifting the covariate density. That study deals with a non-adaptive experiment with covariate choice and regression with importance weighting (Shimodaira, 2000). In contrast, our target is ATE estimation, and we adaptively optimize the covariate density to estimate ATEs. Note that this approach for active learning is an application of importance sampling (Owen, 2013). In causal inference under a covariate shift, Uehara et al. (2020) presents an efficient estimator of policy values, a generalization of ATEs, using double machine learning (DML) (Chernozhukov et al., 2018).
@@
Intuitively, for efficient ATE estimation, we can reduce an estimator of the ATE by using more units for a treatment with a larger variance. In van der Laan (2008), Hahn et al. (2011), and Kato et al. (2020), an experimenter assigns a treatment to an experimental unit with a higher propensity score if it has a larger variance. In this study, we also optimize the choice of experimental units based on their covariates in addition to a propensity score. Under this approach, we use more experimental units if their conditional variances of outcomes are large, conditioned on their covariates. See Figure 2 in Section 4. This covariate choice allows us to reduce an asymptotic variance of an ATE estimator.
The AIPW estimator is popular in adaptive experiments due to its variance reduction and martingale properties. Utilizing these properties, van der Laan (2008), van der Laan & Lendle (2014), Luedtke & van der Laan (2016), and Kato et al. (2020) proposed adaptive experiments for efficiently estimating ATEs. This approach has been further refined by Chernozhukov et al. (2018) as double machine learning with cross-fitting. Kato et al. (2021) refines the results of van der Laan (2008), van der Laan & Lendle (2014), Luedtke & van der Laan (2016), and Chernozhukov et al. (2018), proposing adaptive-fitting. Hadad et al. (2019) and Cook et al. (2023) have focused on this topic, presenting different estimators.
@@
Our framework has a novelty in the optimization of the covariate density, which has not been fully explored in the literature on adaptive experiments for ATE estimation. While active learning focuses on optimizing the covariate density, adaptive experiments typically focus on optimizing the propensity score (Hahn et al., 2011). Following the active learning, we refer to our setting as an active adaptive experiment for ATE estimation.
Contributions: Compared to existing work, our contributions are summarized as follows:
@@ 2.1 Potential Outcomes
Suppose that there are binary treatments $a\in\{1,0\}$. Let us define a corresponding potential outcome by $Y(a)$. Let us also denote a $d$-dimensional covariate by $\mathcal{X}\subset\mathbb{R}^{d}$, where $\mathcal{X}$ is the space. We assume that given $X$, $Y(1)$ and $Y(0)$ are conditionally independent and generated from the conditional densities $r^{1}(y\mid X)$ and $r^{0}(y\mid X)$, respectively.
For any $x\in\mathcal{X}$, let $P_{0}(x)$ be the true conditional distribution of $(Y_{1},Y_{0})$ given $X=x$ whose density is given as $r^{1}(y(1)\mid x)r^{0}(y(0)\mid x)$. Let $\mathbb{E}[\cdot\mid X=x]$ and $\mathrm{Var}\big{(}\cdot\mid X=x\big{)}$ be expectation and variance operators over $P_{0}(x)$. Let us denote the true mean and variance of $Y(a)$ conditioned by $X=x\in\mathcal{X}$ by $\mu_{0}(a)(x)=\mathbb{E}[Y(a)\mid X=x]$ and $\sigma^{2}_{0}(a)(x)=\mathrm{Var}(Y(a)\mid X=x)$, respectively.
@@ 2.2 Average Treatment Effect
For simplicity, $q(x)$ is assumed to be known to the experimenter. This assumption can be mitigated by considering different settings. In Appendix F.1, we introduce another setting where we can observe covariates $X_{t}$ generated from $q(x)$ and decide whether we use it by the rejection sampling.
@@
The covariate density $q(x)$ is known to an experimenter.
@@ 2.3 Active Adaptive Experiments
Consider an adaptive experiment with $T$ rounds, $[T]\coloneqq\{1,2,\dots,T\}$. In the experiment, in each $t\in[T]$, an experimenter decides a covariate density $p_{t}(x)$ and observes an experimental unit with covariates $X_{t}\sim p_{t}(x)$. Given the observed covariates $X_{t}$, the experimenter decides propensity score $w_{t}(a\mid X_{t})$ and assigns treatment $A_{t}\in\{1,0\}$, where $A_{t}\in\{1,0\}$ is an treatment indicator. Then, the experimenter observes an outcome $Y_{t}=\mathbbm{1}[A_{t}=1]Y_{t}(1)+\mathbbm{1}[A_{t}=0]Y_{t}(0)$, where $(Y_{t}(1),Y_{t}(0))$ is an i.i.d. copy of $(Y(1),Y(0))$. After round $T$, the experimenter estimates $\theta_{0}$ by using the observations $\{(Y_{t}(1),Y_{t}(0)\}^{T}_{t=1}$.
@@
Thus, the adaptive experiments yields observations $\{(Y_{t},A_{t},X_{t})\}^{T}_{t=1}$, where $(Y_{t},A_{t},X_{t})$ and $(X_{s},A_{s},Y_{s})$ are correlated over time for $t\neq s$, that is, the samples are not i.i.d. Let $\mathcal{F}_{t-1}=\{X_{t-1},A_{t-1},Y_{t-1},\dots,X_{1},A_{1},Y_{1}\}$ be the history. The probability $w_{t}(a\mid x)$ is a function of a covariate $X_{t}$, an action $A_{t}$, and history $\mathcal{F}_{t-1}$.
@@ 3 Semiparametric Efficiency Bound
This section provides a lower bound for the asymptotic variance of estimators of $\theta_{0}$. We focus on semiparametric efficiency bound for a class of regular estimators. Semiparametric efficiency bounds generalize Cramer-Rao lower bounds, where the latter gives a lower bound for parametric models, while the former gives an efficiency bound for non- and semiparametric models.
@@
$f(y,a,x)=\prod_{d\in\{1,0\}}\Big{\{}r^{d}(y\mid x)w(d\mid x)\Big{\}}^{a=d}p(x).$
Note that in our setting, we can manipulate $w(a\mid x)$ and $p(x)$ and update them during an experiment.
@@
Under this DGP and $q(x)$, we obtain the following theorem. The proof is shown in Appendix B.
@@
Under Assumptions 2.22.2, given $\mathcal{D}_{T}$ and $\widetilde{\mathcal{D}}_{S}$, as $S\to\infty$, the semiparametric efficiency bound for $\theta_{0}$ with fully nonparametric models is
$\tau(w,p)=\mathbb{E}_{X\sim q(x)}\left[\Bigg{\{}\frac{\sigma^{2}_{0}(1)(X)}{w(1\mid X)}+\frac{\sigma^{2}_{0}(0)(X)}{w(0\mid X)}\Bigg{\}}\frac{q(X)}{p(X)}\right].$
@@
This semiparametric efficiency bound can be regarded as a functional of $p$ and $w$ and minimized over them. During our adaptive experiments, we can optimize $p$ and $w$. Note that the semiparametric efficiency gives a lower bound under fixed $p$ and $w$.
@@ 4 Efficient Probabilities
$(w^{*},p^{*}):=\operatorname*{arg\,min}_{(w,p)\in\mathcal{W}\times\mathcal{P}}r(w,p),$ (1)
@@ 4.1 Efficient Propensity Score
First, we find an efficient propensity score that minimizes the semiparametric efficiency bound given some $p(x)$. Given $p$, a stationary point regarding $w$ is given as the point-wise optimizer for each $x\in\mathcal{X}$. That is, given $x$, for $w(x)\coloneqq\{w(a\mid x)\}_{a\in[K]}\in(0,1)^{K}$, we solve $w^{*}(x)=\operatorname*{arg\,max}_{w\in\Delta}\Big{\{}\frac{\sigma^{2}_{0}(1)(x)}{w(1\mid x)}+\frac{\sigma^{2}_{0}(0)(x)}{w(0\mid x)}\Big{\}}\frac{q^{2}(x)}{p(x)}$. Then, for each $x\in\mathcal{X}$, the efficient propensity score is
@@
This efficient propensity score is referred to as the Neyman allocation, which has been traditionally employed in designing experiments with binary treatments without covariate choice (Neyman, 1934; van der Laan, 2008; Hahn et al., 2011; Kato et al., 2020; Adusumilli, 2022).
@@ 4.4 Efficiency Gain
Efficient propensity score.
@@
$\geq\big{(}\sigma_{0}(1)(x)+\sigma_{0}(0)(x)\big{)}^{2}\geq\frac{\sigma^{2}_{0}(1)(x)}{w^{*}(1\mid x)}+\frac{\sigma^{2}_{0}(0)(x)}{w^{*}(0\mid x)}.$
@@
Efficient covariate density.

### v2

@@ Abstract
This study designs an adaptive experiment for efficiently estimating average treatment effects (ATEs). In each round of our adaptive experiment, an experimenter sequentially samples an experimental unit, assigns a treatment, and observes the corresponding outcome immediately. At the end of the experiment, the experimenter estimates an ATE using the gathered samples. The objective is to estimate the ATE with a smaller asymptotic variance. Existing studies have designed experiments that adaptively optimize the propensity score (treatment-assignment probability). As a generalization of such an approach, we propose optimizing the covariate density as well as the propensity score. First, we derive the efficient covariate density and propensity score that minimize the semiparametric efficiency bound and find that optimizing both covariate density and propensity score minimizes the semiparametric efficiency bound more effectively than optimizing only the propensity score. Next, we design an adaptive experiment using the efficient covariate density and propensity score sequentially estimated during the experiment. Lastly, we propose an ATE estimator whose asymptotic variance aligns with the minimized semiparametric efficiency bound.
Treatment effect estimation, Adaptive experimental design, Active learning, Causal inference
@@ 1 Introduction
Experimental approaches play a pivotal role in uncovering causality in various fields of science and industrial applications, such as epidemiology and economics (Shein-Chung Chow, 2005; Hahn et al., 2011; Chow & Chang, 2011; FDA, 2019). This study focuses on cases with binary treatments. Typically, one of these treatments corresponds to the treatment, and the other to the control (Imbens & Rubin, 2015). Among the quantities representing causal effects, this study focuses on an average treatment effect (ATE), which defines a causal effect as the difference between the expected outcomes of binary treatments.
Treatments are an abstraction of an experimenter’s choices, including drugs, online advertisements, and economic policies, which are also referred to differently, such as actions and arms. A difference in outcomes of binary treatments (treatment effect) is often considered a causal effect (Imbens & Rubin, 2015). However, due to the counterfactual property, individual treatment effects cannot be observed. Therefore, we use ATEs to capture causality.
A fundamental experimental approach for estimating ATEs is a randomized control trial (RCT). In an RCT, one of the binary treatments is randomly assigned to each experimental unit to obtain an unbiased estimator of ATEs. While RCTs are the gold standard and reliable (Hariton & Locascio, 2018), they often require large sample sizes, which can be costly. To reduce the sample size as much as possible, or to reduce an estimation error given a sample size, adaptive experimental designs have gained increasing attention across various fields (van der Laan, 2008; Hahn et al., 2011; Chow & Chang, 2011; Villar et al., 2015; FDA, 2019).
For estimating ATEs efficiently (with smaller asymptotic variances), van der Laan (2008) and Hahn et al. (2011) design adaptive experiments that optimize the propensity score (treatment-assignment probability). They show that by optimizing the propensity score, we can minimize the asymptotic variance of ATE estimators. Note that an asymptotic variance can be interpreted as an asymptotic mean squared error (MSE) of ATE estimation. Furthermore, in hypothesis testing, as an asymptotic variance decreases, the sample size required for hypothesis testing also decreases. Tabord-Meehan (2022), Kato et al. (2020), and Cook et al. (2023) design refined experiments based on this approach.
@@
Firstly, we derive the semiparametric efficiency bound, a lower bound of the asymptotic variance, which is a functional of a covariate density and a propensity score. Based on this efficiency bound, we compute the efficient covariate density and propensity score that minimize the lower bound. Then, we design an adaptive experiment that sequentially estimates the efficient covariate density and propensity score. In the proposed experiment, an experimental unit is sampled following an adaptively estimated efficient covariate density and randomly assigned treatments following an estimated efficient propensity score. At the experiment’s end, the experimenter obtains an estimator whose asymptotic variance aligns with the semiparametric efficiency bound. We employ the Augmented Inverse Probability Weighting (AIPW) estimator with covariate shift adaptation using Importance Weighting (AIPWIW). We refer to the setting of an adaptive experiment as active adaptive experimental design with covariate choice and our designed experiment as the Active-Adaptive-Sampling (AAS)-AIPWIW experiment.
@@
Our method is inspired by three lines of work: adaptive experimental design for efficient ATE estimation, active learning using a covariate shift, and off-policy evaluation under a covariate shift. van der Laan (2008), Hahn et al. (2011), Tabord-Meehan (2022), and Kato et al. (2020) design adaptive experiments for estimating ATEs with smaller asymptotic variances by optimizing propensity scores. Our result generalizes their work by incorporating covariate density optimization. In active learning of regression models, Sugiyama (2006) demonstrates that the asymptotic variance of regression coefficient estimators can be reduced by appropriately shifting the covariate density. That study deals with a non-adaptive experiment and regression with importance weighting (Shimodaira, 2000). In contrast, our target is ATE estimation, and we adaptively optimize the covariate density to estimate ATEs. Note that this approach for active learning is an application of importance sampling (Owen, 2013). In causal inference under a covariate shift, Uehara et al. (2020) presents an efficient estimator of policy values, a generalization of ATEs, using double machine learning (DML, Chernozhukov et al., 2018).
@@
Intuitively, we can reduce the asymptotic variance of the ATE estimator by using more units for a treatment with a larger variance. In existing studies, an experimenter assigns a treatment to an experimental unit with a higher propensity score if it has a larger variance. In this study, we also optimize the choice of experimental units based on their covariates in addition to the propensity score. Under this approach, we use more experimental units if their conditional variances of outcomes are large, conditioned on their covariates. See Figure 2 in Section 4. This covariate choice allows us to reduce the asymptotic variance of an ATE estimator.
@@
Our framework introduces a novel aspect of optimizing covariate density, which has not been fully explored in the literature on adaptive experiments for ATE estimation. While active learning focuses on optimizing the covariate density, adaptive experiments typically focus on optimizing the propensity score. We refer to our setting as an active adaptive experiment for ATE estimation.
Contributions
Compared to existing work, our contributions are summarized as follows:
@@ 2.1 Potential Outcomes
There is a binary treatment $a\in\{1,0\}$. Let us define the corresponding potential outcome by $Y(a)$. Let $X\in\mathcal{X}\subset\mathbb{R}^{d}$ be a $d$-dimensional covariate, where $\mathcal{X}$ is the space.
For any $x\in\mathcal{X}$, let $P_{0}(x)$ be the true conditional distribution of $(Y_{1},Y_{0})$ given $X=x$ whose density is given as $r^{1}(y(1),y(0)\mid x)$. Let $\mathbb{E}[\cdot\mid X=x]$ and $\mathrm{Var}\big{(}\cdot\mid X=x\big{)}$ be expectation and variance operators over $P_{0}(x)$. Let us denote the true mean and variance of $Y(a)$ conditioned by $X=x\in\mathcal{X}$ by $\mu_{0}(a)(x)=\mathbb{E}[Y(a)\mid X=x]$ and $\sigma^{2}_{0}(a)(x)=\mathrm{Var}(Y(a)\mid X=x)$, respectively.
@@ 2.2 Average Treatment Effect
For simplicity, $q(x)$ is assumed to be known to the experimenter. This assumption can be mitigated by considering different settings. In Appendix E.1, we introduce another setting where we can observe covariates $X_{t}$ generated from $q(x)$ and decide whether we use it by the rejection sampling.
@@
The covariate density $q(x)$ is known.
@@ 2.3 Active Adaptive Experiments
Consider an adaptive experiment with $T$ rounds, $[T]\coloneqq\{1,2,\dots,T\}$. In the experiment, in each $t\in[T]$, an experimenter decides a covariate density $p_{t}(x)$ and observes an experimental unit with covariates $X_{t}\sim p_{t}(x)$. Given the observed covariates $X_{t}$, the experimenter decides propensity score $w_{t}(a\mid X_{t})$ and assigns treatment $A_{t}\in\{1,0\}$, where $A_{t}\in\{1,0\}$ is a treatment indicator. Then, the experimenter observes an outcome $Y_{t}=\mathbbm{1}[A_{t}=1]Y_{t}(1)+\mathbbm{1}[A_{t}=0]Y_{t}(0)$, where $(Y_{t}(1),Y_{t}(0))$ is an i.i.d. copy of $(Y(1),Y(0))$. After round $T$, the experimenter estimates $\theta_{0}$ by using the observations $\{(Y_{t}(1),Y_{t}(0)\}^{T}_{t=1}$.
@@
Thus, the adaptive experiments yields observations $\{(Y_{t},A_{t},X_{t})\}^{T}_{t=1}$, where $(Y_{t},A_{t},X_{t})$ and $(Y_{s},A_{s},X_{s})$ are correlated over time for $t\neq s$, that is, the samples are not i.i.d. Let $\mathcal{F}_{t-1}=\{X_{t-1},A_{t-1},Y_{t-1},\dots,X_{1},A_{1},Y_{1}\}$ be the history. The probability $w_{t}(a\mid x)$ is a function of a covariate $X_{t}$, an action $A_{t}$, and history $\mathcal{F}_{t-1}$.
@@ 3 Semiparametric Efficiency Bound
This section provides a lower bound for the asymptotic variance of ATE estimators. We focus on semiparametric efficiency bound for a class of regular estimators. The semiparametric efficiency bound generalize the Cramer-Rao lower bound, where the latter considers parametric models, while the former considers non- and semiparametric models.
@@
$f(y,a,x)=\prod_{d\in\{1,0\}}\left(r^{d}(y(d)\mid x)w(d\mid x)\right)^{a=d}p(x),$
where $r^{d}(y(d)\mid x)$ is the marginal distribution of $Y(d)$. Note that in our setting, we can manipulate $w(a\mid x)$ and $p(x)$ and update them during an experiment.
@@
Under this DGP and $q(x)$, we obtain the following theorem. The proof is shown in Appendix A.
@@
Under Assumption 2.2, given $\mathcal{D}_{T}$ and $\widetilde{\mathcal{D}}_{S}$, as $S\to\infty$, the semiparametric efficiency bound for $\theta_{0}$ with fully nonparametric models is
$\tau(w,p)=\mathbb{E}_{X\sim q(x)}\left[\left(\frac{\sigma^{2}_{0}(1)(X)}{w(1\mid X)}+\frac{\sigma^{2}_{0}(0)(X)}{w(0\mid X)}\right)\frac{q(X)}{p(X)}\right].$
@@
This semiparametric efficiency bound can be regarded as a functional of $p$ and $w$. During our adaptive experiments, we can minimize the bound regarding $p$ and $w$.
@@ 4 Efficient Probabilities
$(w^{*},p^{*}):=\operatorname*{arg\,min}_{(w,p)\in\mathcal{W}\times\mathcal{P}}\tau(w,p),$ (1)
@@ 4.1 Efficient Propensity Score
First, we find an efficient propensity score that minimizes the semiparametric efficiency bound given some $p(x)$. Given $p$, a stationary point regarding $w$ is given as the point-wise optimizer for each $x\in\mathcal{X}$. That is, given $x$, for $w(x)\coloneqq\{w(a\mid x)\}_{a\in[K]}\in(0,1)^{K}$, we have $w^{*}(x)=\operatorname*{arg\,max}_{w\in\Delta}\Big{\{}\frac{\sigma^{2}_{0}(1)(x)}{w(1\mid x)}+\frac{\sigma^{2}_{0}(0)(x)}{w(0\mid x)}\Big{\}}\frac{q^{2}(x)}{p(x)}$. Then, for each $x\in\mathcal{X}$, the efficient propensity score is
@@
This efficient propensity score is referred to as the Neyman allocation, which has been traditionally employed in designing experiments with binary treatments without covariate choice (Neyman, 1934; van der Laan, 2008; Hahn et al., 2011; Kato et al., 2020; Adusumilli, 2022; Dai et al., 2023).
@@ 4.4 Efficiency Gain
Efficient propensity score
@@
$\geq\big{(}\sigma_{0}(1)(x)+\sigma_{0}(0)(x)\big{)}^{2}=\frac{\sigma^{2}_{0}(1)(x)}{w^{*}(1\mid x)}+\frac{\sigma^{2}_{0}(0)(x)}{w^{*}(0\mid x)}.$
@@
Efficient covariate density